# DABL: Detecting Semantic Anomalies in Business Processes Using Large Language Models

### v1

@@ Introduction
Interpretation of the Cause of Anomalies
Table 3: The results of DABL in interpreting the causes of anomalies.
@@
Table 4: Irregularity patterns identified in the travel permit log.
@@
ID Example Trace Output Cause
A1 [CF, SF, IFN, AP, P, P, P, P] The activities ’Payment’ and ’Payment’ are reworked after ’Payment’.
A2 [CF, P, P, SF, IFN, AP, P]
A3 [CF, SF, IFN, RRAP, IDAP, SAP, AP, P]
The activity ’Receive Result Appeal from Prefecture’ is executed too early, it should be executed after ’Send Appeal to Prefecture’.
Table 3 shows the results of DABL in interpreting the causes of anomalies. On dataset $\mathcal{D}_{1}$, both ROUGE-2 and ROUGE-L scores are relatively high, indicating that DABL performs well in identifying the causes of anomalies, even for processes not included in the training data. For dataset $\mathcal{D}_{2}$, DABL exhibits better performance because the normal behaviors of the processes are well-represented in the training data. Furthermore, the slight difference between recall and precision suggests that the model maintains a good balance. These results demonstrate that DABL is effective at interpreting the causes of anomalies in terms of both bigrams and longest common subsequences.
However, these result values may be underestimated because the cause of an anomaly can be interpreted in various ways. For example, for a desired trace [A, B, C, D, E], the anomaly [A, B, E, C, D] can be interpreted as ”The activities ’C’ and ’D’ are executed too late, they should be executed after ’B’” and ”The activity ’E’ is executed too early, it should be executed after ’D’.” Nevertheless, we only provide one reference answer to calculate ROUGE-2 and ROUGE-L scores.
Qualitative Examples
Figure 5: Qualitative results of various LLMs on an early anomaly.
Fig. 5 and Fig. 6 illustrate the performance of existing LLMs and DABL on early and exclusion anomalies, respectively. Due to the absence of business process-related knowledge, existing LLMs may not offer definitive answers. In contrast, our DABL is proficient in identifying anomalies and interpreting their underlying causes, thanks to the integration of a vast array of process models from diverse domains in the training data.
@@
We apply our DABL on a real-world permit log from the BPI 2020 challenge (van Dongen 2020), which captures data on work trips conducted by university employees. The process flow involves the request for and approval of a travel permit, the trip itself, a subsequent travel declaration, as well as associated reimbursements.
@@
We apply our DABL on another real-world event log from an information system managing road traffic fines (de Leoni and Mannhardt 2015), which captures the road traffic fine management process. The process flow involves the creation of a fine, appeal to the prefecture, addition of penalties, and fine payment.
This log contains 150,370 traces with 231 variants. DABL identifies 56 anomalous variants. These detected anomalies reveal some interesting irregularity patterns as detailed in Table 5. The examples illustrate irregularity patterns where the fine is repeatedly paid (A1), the fine is paid before the penalty is added (A2), and the result appeal from the prefecture is received before the appeal is sent to the prefecture (A3).
@@ Acknowledgments
This work is supported by China National Science Foundation (Granted No. 62072301).

### v2

@@ Introduction
Compared Methods
In our evaluation, we compare our DABL to existing semantic business process anomaly detection methods: SENSE (caspary2023does) and SEM (van2021natural). SENSE offers both SVM-based and BERT-based models for detecting anomalous event pairs, which we denote as SENSE-SVM and SENSE-BERT, respectively. These methods divide traces into event pairs and determine whether each pair is normal or anomalous. If at least one event pair in a trace is identified as anomalous, the entire trace is classified as anomalous. It is important to note that SEM can only detect anomalous event pairs that share the same business object, automatically classifying pairs with distinct business objects as non-anomalous. Due to the high training costs, we utilize the open-source trained models provided by the authors for dataset $\mathcal{D}_{1}$. For dataset $\mathcal{D}_{2}$, we train the comparative models using the 21,298 normal traces available within it. The hyper-parameters of these methods are set to the values that yielded the best results reported in the original paper.
Evaluation Metrics
Following existing anomaly detection methods, we employ precision, recall, F1-score and accuracy to evaluate the anomaly detection performance. The recall-oriented understudy for gisting evaluation (ROUGE) (lin2004rouge) is a software package and metric set designed to assess the quality of generated text by comparing it with ground truth text. In our evaluation of DABL’s ability to interpret the cause of anomalies, we utilize ROUGE-2 and ROUGE-L metrics. We conduct each experiment five times and report the mean results.
@@
Table 3: The results of DABL in interpreting the causes of anomalies.
Quantitative Results
Anomaly Detection
Figure 5: Qualitative results of various LLMs on an early anomaly.
Interpretation of the Cause of Anomalies
Table 3 shows the results of DABL in interpreting anomalies’ causes. On dataset $\mathcal{D}_{1}$, both ROUGE-2 and ROUGE-L scores are relatively high, indicating that DABL performs well in identifying anomalies’ causes, even for processes not included in the training data. For dataset $\mathcal{D}_{2}$, DABL exhibits better performance because the normal behaviors of the processes are well-represented in the training data. Furthermore, the slight difference between recall and precision suggests that the model maintains a good balance. These results demonstrate that DABL is effective at interpreting anomalies’ causes regarding both bigrams and longest common subsequences.
[…]
Qualitative Examples
Fig. 5 and Fig. 6 illustrate the performance of existing LLMs and DABL on early and exclusion anomalies, respectively. Due to the absence of business process-related knowledge, existing LLMs may not offer definitive answers. In contrast, DABL is proficient in identifying anomalies and interpreting their causes, thanks to the integration of a vast array of process models from diverse domains in the training data.
@@
Table 4: Irregularity patterns identified in the travel permit log.
ID Example Trace Output Cause
[…]
The activity ’Receive Result Appeal from Prefecture’ is executed too early,
it should be executed after ’Send Appeal to Prefecture’.
@@
We apply our DABL on a real-world travel permit log from the BPI 2020 challenge (BPIC20), which captures data on work trips conducted by university employees. The process flow involves the request for and approval of a travel permit, the trip itself, a subsequent travel declaration, as well as associated reimbursements.
@@
We apply our DABL on another real-world event log from an information system managing road traffic fines (RTFMP), which captures the road traffic fine management process. The process flow involves the creation of a fine, appeal to the prefecture, addition of penalties, and fine payment.
This log contains 150,370 traces with 231 variants. DABL identifies 56 anomalous variants, which reveal some interesting irregularity patterns as detailed in Table 5. The examples illustrate irregularity patterns where the fine is repeatedly paid (A1), the fine is paid before the penalty is added (A2), and the result appeal from the prefecture is received before the appeal is sent to the prefecture (A3).
Limitations
Despite collecting a significant amount of data to fine-tune LLMs for detecting business process anomalies and achieving good performance, it is important to acknowledge some limitations of DABL.
First, the precision of the collected process model may not be equal to 1, which means it might allow traces not observed in the original process. This discrepancy could impact the quality of the training data, posing a limitation for DABL. Second, our open-source model may struggle with traces that, while irregular from a common-sense perspective, are normal within their customized processes. Finally, when multiple anomaly types occur within a trace, DABL can still classify the trace as an anomaly but can only interpret one type of anomaly at a time. After fixing the interpreted anomaly and re-inputting the trace into DABL, the model will identify and interpret another anomaly type.
@@ Acknowledgments
This work is supported by China National Science Foundation (Granted Number 62072301). This work is also partially supported by the Program of Technology Innovation of the Science and Technology Commission of Shanghai Municipality (Granted No.22DZ1100103).